Operations Report – January 22nd

Crew 220 Operations Report

SOL: 3

Submitting Report: Lee Roberts, Crew Engineer

Non-nominal systems: None

Generator: Standby Rental Generator

Hours Run: 0

Generator running times

Start Time: N/A

End Time: N/A

Total Run Time: 0 hours

Solar SOC: 100%

Diesel Reading: 49%

Station Propane Reading: 24%

Ethanol Free Gasoline: N/A

Water in Loft Tank: 30 Gallons

Water Meter: 148045.7

Water in Static Tank: 395 Gallons

Static to Loft Tank: 49 Gallons

Water in GreenHab Tank: 222 Gallons

Water in Science Dome: N/A

Toilet Tank Emptied: No

Spirit Rover Used: No

Ending Hours: 127.7

Beginning Charge: N/A

Ending Charge: 100%

Curiosity Rover Used: No

Ending Hours: 133.7

Beginning Charge: N/A

Ending Charge: 100%

Notes on Rovers: None

ATVs Used: None

HabCar Used, Why, Where? Negative

Crew Car Used, Why, Where? Negative

General Notes and Comments: None

Summary of Internet: While restoring Hab power the Internet router failed to power up properly. After several attempts to reboot the router, Outpost was contacted and advised of the concern. The router was sent to them, and they are working on repairing or replacing it. Two crew members from the MAU team brought a router that they had in their possession and installed it as a temporary replacement, with the permission of Outpost.

Summary of Suits and Radios: No EVAs were conducted from MDRS today, but crew members drilled in donning and doffing procedures. Suits and radios are all charging after restoration of power, and all systems appear to be operating nominally.

Summary of Hab Operations: We were still in emergency operations for the first portion of the day, but power was restored with a standby rental generator at approximately 0900. The decision was made to restore power to electrical systems with a staggered ramp up, beginning with radio comms, then suits, rovers, and Internet, and then full load after the SOC was stable above 50%. All systems appear to be operating nominally at this time. Our crew is extremely grateful for the dedicated effort of the Outpost team in restoring power for us.

Summary of GreenHab Operations: Please refer to the GreenHab Report

Summary of RAM Operations: No operations today

Summary of Observatory Issues: None

Summary of Health and Safety Issues: Please Refer to the Health and Safety Report

Questions, concerns, and requests to Mission Support:We have no specific concerns or requests at this time.

Copyright © The Mars Society. All rights reserved. | Main Site